OpenStudy (anonymous):

How do you factor 5c^3+2c^2-5cd^2-2d^2????? 😔😔😔😔😔😔

OpenStudy (perl):

did you try grouping

OpenStudy (anonymous):

No I haven't

pooja195 (pooja195):

I dont think this is something that can be factored :/

pooja195 (pooja195):

\[\huge~5c^3−5cd^2+2c^2−2d^2 \] tthis is all you would get just organized .-.

pooja195 (pooja195):

no like terms either :/

OpenStudy (anonymous):

Thx u

pooja195 (pooja195):

id if i helped :/

OpenStudy (anonymous):

Now I'm stuck lol

OpenStudy (perl):

This does factor. \[\Large { 5c^3+2c^2-5cd^2-2d^2 \\= c^2 (5c + 2) - d^2( 5c + 2) \\ = (5c+2)( c^2 - d^2) \\ = (5c+2)( c -d )( c + d) }\]

pooja195 (pooja195):

5cd^2??

OpenStudy (perl):

you can also factor that way if you like

OpenStudy (perl):

\[\huge{5c^3−5cd^2+2c^2−2d^2 \\ =5c ~(c^2 - d^2) + 2( c^2 - d^2) \\ = (c^2 - d^2) ( 5c+2) \\= (c-d)(c+d)(5c+2) } \]

pooja195 (pooja195):

good job perl :)

OpenStudy (anonymous):

Thx u

OpenStudy (perl):

Thanks :)